OpenStudy (anonymous):

does any body no this evulate the expression when p=-24 and q=4 2p -q

OpenStudy (anonymous):

First, substitute into the expression 2p - q = 2(-24) - (4) = -48 - 4 = -52

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np

OpenStudy (anonymous):

do you no 39615.10638 rounded to the nearest hundredths place

OpenStudy (anonymous):

39615.11